OpenStudy (anonymous):

When 3 coins are tossed, what is the probability of getting (i) 2 heads and a tail (ii) at least 2 tails

OpenStudy (kropot72):

There are 8 possible outcomes when 3 fair coins are tossed: TTT TTH THT THH HTT HTH HHT HHH How many of the outcomes have 2 heads and a tail?

OpenStudy (anonymous):

Notice that saying "probability of 2 heads and a tail" is the same as saying "the probability of 3 heads", since the chance of getting a tail or a heads is equal, i.e 1/2. Hence, to get the probability of 3 heads on 3 coins, it's just 1/2 x 1/2 x 1/2 = ?

OpenStudy (kropot72):

If the binomial distribution is used to solve (i) the correct calculation is found as follows: \[P(2\ heads)=3C2\times 0.5^{2}\times 0.5=3C2\times 0.5^{3}\]

OpenStudy (anonymous):

The answer for it is (i) 1/72 (ii) 1/54 but i do not know how to arrive at it

OpenStudy (anonymous):

no it isn't two heads and a tail THH HTH HHT three out of eight

OpenStudy (kropot72):

@amstro Are you sure that you have posted the question correctly. The answer that you have posted for (i) is definitely not correct for the question as posted.

OpenStudy (anonymous):

@kropot72 i checked the question it is correct.

OpenStudy (kropot72):

Then the answer that you posted is incorrect. The correct answer has been confirmed by @satellite73